# Synthetic Socratic Debates: Examining Persona Effects on Moral Decision and Persuasion Dynamics

### v1

@@ 5 Discussions
5.1 Analysis of Moral Judgments
@@
We find that GPT-4o’s no-persona responses yield higher moral scores than all its persona-conditioned counterparts, suggesting that assigned personas systematically bias the model toward greater blame of the narrator. In contrast, LLaMA 4 exhibits lower scores in the no-persona condition compared to its persona-assigned outputs. These findings suggest that moral judgments in large language models are not neutral by default and that assigning personas introduces systematic variation in moral evaluations.
5.2 Analyses on debate dynamics
@@
The results presented in Appendix L show a consistent increase in confidence across all persona dimensions and groups by the end of the debate. Interestingly, this increase occurs regardless of whether the agent’s choice changes during the debate. These findings suggest that agent responses become more robust over time, and that this growing trend is largely independent of persona characteristics.
@@
Unlike confidence scores, mode of persuasion scores exhibit a different trend. Using the same analytical approach as in confidence analysis, we report the results in Appendix M. We find that scores for all three modes of persuasion, Pathos, Ethos, and Logos, generally decline over the course of the debate, regardless of persona characteristics. Among them, Pathos shows the largest decrease, while Logos shows the least. With the exception of the Big Five personality dimension, nearly all of these decreases are statistically significant under a one-sided $t$-test at the 0.05 significance level. This suggests that as debates progress, models tend to adopt weaker persuasive strategies across personas.
@@ 6 Conclusion
Prompt-constructed personas inherit biases from both the language model’s pre-training data and the researchers’ design choices. Some persona combinations may inadvertently encode harmful stereotypes (e.g. linking political ideology with moral rigidity). Because these personas guide the model’s argumentative stance, they can amplify or legitimize biased moral frameworks. To minimize harm, we manually reviewed prompt templates for discriminatory language, disallowed protected-attribute slurs, and released all templates under a harm-reporting protocol so that stakeholders can flag problematic content.
@@
Persona-conditioned moral debate systems could be used to sway public opinion or fabricate grassroots consensus by selectively deploying persuasive personas. Malicious actors might exploit high-influence personas to shape moral discourse on sensitive topics such as elections or public health. Although our study is purely analytical, we advocate for guardrails, such as transparent persona disclosure, provenance tracking, and rate limiting, to deter covert mass persuasion. We also encourage policymakers to adopt audit requirements for the large-scale deployment of persuasive conversational agents.
@@
All data and models used in this work are covered under academic licenses permitting research use. We strictly adhere to the intended use policies and do not involve any sensitive or personally identifiable data. Upon acceptance, we plan to open-source the dataset for research purposes. The models employed were accessed via commercial APIs, with a total usage cost of \$500.
@@ Appendix B Persona Modeling Details
These strings are embedded directly into the prompts used for persona construction and model input.
@@ Appendix D Multi-Turn Debate Details
We select personas based on the results from RQ1. Specifically, for each moral dilemma scenario, we identify all persona pairs that express divergent moral judgments. From these, we randomly sample 10 persona pairs per scenario for use in the debate simulations.
@@ Appendix F Moral Foundation Theory Details
MFT posits that human moral reasoning is organized around a small set of evolutionarily ancient, culturally elaborated “foundations.” Originally five in number—Care/Harm, Fairness/Cheating, Loyalty/Betrayal, Authority/Subversion, and Sanctity/Degradation—these dimensions were distilled from cross-cultural anthropology, comparative psychology, and large-scale survey factor analyses Haidt and Graham (2007); Graham et al. (2013). Each foundation functions as an intuitive template for evaluating social actions and institutions, thereby explaining systematic differences in political and cultural moral judgments.
@@
Subsequent work showed that a distinct Liberty/Oppression foundation (capturing concerns about personal autonomy and resistance to coercion) consistently emerges as an independent factor in exploratory and confirmatory analyses of Moral Foundations Questionnaire items Iyer et al. (2012). Adopting the six-factor variant improves (i) construct coverage, by recognizing libertarian moral intuitions overlooked in the five-factor model; (ii) predictive validity, yielding finer-grained correlations with political ideology and policy attitudes; and (iii) cross-cultural robustness, as Liberty loads separately in diverse national samples. Consequently, we model moral preferences along these six orthogonal axes to capture a broader spectrum of value conflict in persona-conditioned debates.
@@ Appendix G Persuasion Effectiveness Details
We evaluate persuasion effectiveness using four metrics: stance shifts, win rate, consensus rate, and efficiency, each grounded in theoretical and empirical motivations. Stance shifts capture individual belief changes, a core feature in classical and computational rhetoric (Tan et al., 2016). Win rate reflects success against an opposing view, aligning with debate contexts where the goal involves persuasion in addition to reasoning. Consensus rate measures cooperative outcomes, indicating whether moral convergence emerges, a key aim in deliberation and democratic theory (Mercier and Sperber, 2011; Johnson, 1991). Efficiency assesses the cost of achieving agreement.
@@ I.1 Moral Foundation Theory Results
Figure 6 illustrates how different persona groups engage with moral foundation dimensions in their moral judgments.
@@ I.2 Mode of Persuasion Results
Figure 7 illustrates how different persona groups employ various modes of persuasion in their debate process.
(a) age
(b) gender
(c) country
(d) social class
(e) ideology
(f) big5
Figure 7: persona impact on persuasion modes
Appendix J Additional Results for LLaMA-4-Maverick
J.1 Moral Foundation Theory Results
Figure 8 illustrates how different persona groups engage with moral foundation dimensions in their moral judgments.
(a) age
(b) gender
(c) country
(d) social class
(e) ideology
(f) big5
J.2 Quantifying Moral Judgment Results
Figure 9 presents the moral judgment scores of LLaMA-4-Maverick.
J.3 Impact of Persona on Persuasion Metrics
Figure 10 presents the persona impact on persuasion effectiveness metrics for LLaMA-4-Maverick.
@@
Figure 10: Persona impact on persuasion effectiveness for LLaMA-4-Maverick, measured by consensus ratio and win ratio.
Appendix K Debate Pool Statistics
Table 1 to Table 6 present the distributions of personas across different dimensions within the entire debate pool.
@@
Appendix L Debate Dynamics: Confidence Changes
Table 7 to Table 12 present the changes in model confidence and their statistical significance, as measured by $t$-tests, across different persona dimensions.
@@
40 $-0.27$ $-0.01$ $5.44\times 10^{-35}$ Yes
50 $-0.26$ $-0.01$ $6.92\times 10^{-33}$ Yes
30 $-0.28$ $-0.01$ $1.21\times 10^{-48}$ Yes
20 $-0.30$ $-0.01$ $9.32\times 10^{-38}$ Yes
60 $-0.28$ $-0.01$ $1.30\times 10^{-46}$ Yes
@@
Non-binary $-0.29$ $-0.01$ $6.16\times 10^{-74}$ Yes
Male $-0.26$ $-0.01$ $2.05\times 10^{-63}$ Yes
Female $-0.29$ $-0.01$ $3.11\times 10^{-60}$ Yes
@@
India $-0.26$ $-0.01$ $4.06\times 10^{-37}$ Yes
Brazil $-0.26$ $-0.01$ $3.85\times 10^{-33}$ Yes
China $-0.25$ $-0.01$ $6.80\times 10^{-32}$ Yes
Nigeria $-0.27$ $-0.01$ $6.06\times 10^{-37}$ Yes
France $-0.36$ $-0.01$ $9.63\times 10^{-34}$ Yes
United States $-0.28$ $-0.01$ $2.60\times 10^{-32}$ Yes
@@
Middle class $-0.30$ $-0.01$ $7.06\times 10^{-73}$ Yes
Upper class $-0.26$ $-0.01$ $1.25\times 10^{-58}$ Yes
Lower class $-0.28$ $-0.01$ $4.33\times 10^{-66}$ Yes
@@
Libertarian-Right $-0.29$ $-0.01$ $5.20\times 10^{-53}$ Yes
Libertarian-Left $-0.30$ $-0.01$ $2.32\times 10^{-60}$ Yes
Authoritarian-Left $-0.30$ $-0.01$ $1.26\times 10^{-44}$ Yes
Authoritarian-Right $-0.22$ $-0.01$ $3.45\times 10^{-42}$ Yes
@@
High Extraversion $-0.30$ $-0.01$ $2.38\times 10^{-24}$ Yes
High Conscientiousness $-0.28$ $-0.01$ $4.88\times 10^{-22}$ Yes
Low Conscientiousness $-0.23$ $-0.02$ $3.84\times 10^{-14}$ Yes
High Neuroticism $-0.25$ $-0.01$ $3.34\times 10^{-20}$ Yes
High Openness $-0.28$ $-0.01$ $2.10\times 10^{-16}$ Yes
Low Extraversion $-0.29$ $-0.01$ $5.65\times 10^{-16}$ Yes
Low Agreeableness $-0.35$ $-0.00$ $1.38\times 10^{-44}$ Yes
Low Neuroticism $-0.26$ $-0.02$ $2.62\times 10^{-21}$ Yes
High Agreeableness $-0.33$ $-0.01$ $2.90\times 10^{-18}$ Yes
Low Openness $-0.19$ $-0.01$ $1.86\times 10^{-13}$ Yes
@@
Appendix M Debate Dynamics: Mode of Persuasion Score Changes
Table 13 to Table 18 present the changes in model’s mode of persuasions and their statistical significance, as measured by $t$-tests, across different persona dimensions.

### v2

@@ 5 Discussions
5.1 Analyses of Moral Judgments
@@
We find that GPT-4o’s no-persona responses yield higher moral scores than all its persona-conditioned counterparts, suggesting that assigned personas systematically bias the model toward greater blame of the narrator. In contrast, LLaMA-4-Maverick and Qwen3-235B-A22B exhibit lower scores in the no-persona condition compared to its persona-assigned outputs. These findings suggest that moral judgments in large language models are not neutral by default and that assigning personas introduces systematic variation in moral evaluations.
Third-Person v.s. First-Person Perspective
Although we lack human-annotated data for direct comparison, we experimentally examine how LLMs behave when prompted from a first-person perspective. We hypothesize that when the model is instructed to respond as one of the parties in a dilemma, it adopts that character’s viewpoint and tends to shift blame onto the opposing party. To test this, we construct two first-person prompts: first-person narrator (“Imagine that you are the narrator of the following moral dilemma”) and first-person opponent (“Imagine that you are the individual mentioned in the following moral dilemma, opposed to the narrator”). We randomly select 100 personas and compute their average moral judgment scores (Table˜1). The results support our hypothesis that adopting a first-person perspective substantially shapes the model’s moral judgments.
Perspective GPT-4o LLaMA-4-Maverick
Third-person 2.72 $\pm$ 1.19 2.71 $\pm$ 1.26
First-person (narrator) 3.30 $\pm$ 1.48 3.83 $\pm$ 1.38
First-person (opponent) 1.82 $\pm$ 1.13 2.28 $\pm$ 1.25
Table 1: Moral judgment scores across different prompting perspectives for GPT-4o and LLaMA-4-Maverick.
Is There Any Positional Bias?
To mitigate positional bias, we use Likert-scale ratings rather than binary choice questions. To test robustness, we reversed the order of the two actions and evaluated the models’ responses. For GPT-4o, the average score shifted only slightly (2.72 to 2.59), remaining in the author-blameworthy range and confirming the consistency of our findings. In contrast, LLaMA-4-Maverick’s average score shifted from 2.71 to 3.05, moving into the neutral range. These results suggest that GPT-4o is relatively robust to action-order changes, whereas LLaMA-4-Maverick is more sensitive to positional effects.
5.2 Analyses of Debate Dynamics
@@
The results presented in Appendix˜N show a consistent increase in confidence across all persona dimensions and groups by the end of the debate. Interestingly, this increase occurs regardless of whether the agent’s choice changes during the debate. These findings suggest that agent responses become more robust over time, and that this growing trend is largely independent of persona characteristics.
@@
Unlike confidence scores, mode of persuasion scores exhibit a different trend. Using the same analytical approach as in confidence analysis, we report the results in Appendix˜O. We find that scores for all three modes of persuasion, Pathos, Ethos, and Logos, generally decline over the course of the debate, regardless of persona characteristics. Among them, Pathos shows the largest decrease, while Logos shows the least. With the exception of the Big Five personality dimension, nearly all of these decreases are statistically significant under a one-sided $t$-test at the 0.05 significance level. This suggests that as debates progress, models tend to adopt weaker persuasive strategies across personas.
Does Debate Order Affect Model Behavior?
We investigate whether reversing the debate order for each pair of personas influences their confidence scores or the overall consensus ratio. To test this, we conducted two-sided t-tests on these outcomes. The results in Appendix˜P show no significant differences across metrics or models (all $p>0.17$). Thus, debate order does not significantly impact model behavior in terms of confidence or consensus—particularly for GPT-4o, which appears highly robust to this variation.
5.3 Cross-Model Differences
Overall, the findings are broadly consistent across models for persona dimensions that show statistically significant effects, including both moral judgment score distributions and debate dynamics. However, Qwen3-235B-A22B stands out with a notably different pattern. Unlike the other models, which consistently exhibit a bias toward blaming the narrator regardless of persona assignment, Qwen3 more often shifts blame to others. These differences suggest that, while general trends are shared, individual models may encode distinct biases. This highlights the importance of including multiple LLMs in such analyses to avoid drawing conclusions that reflect idiosyncrasies of a single model rather than robust, generalizable patterns (Chakraborty et al., 2025; Nabizadeh et al., 2025).
@@ 6 Conclusion
Prompt‐constructed personas inherit biases from both the language model’s pre‐training data and the researchers’ design choices. Some persona combinations may inadvertently encode harmful stereotypes (e.g. linking political ideology with moral rigidity). Because these personas guide the model’s argumentative stance, they can amplify or legitimize biased moral frameworks. To minimize harm, we manually reviewed prompt templates for discriminatory language, disallowed protected‐attribute slurs, and released all templates under a harm‐reporting protocol so that stakeholders can flag problematic content.
@@
Persona‐conditioned moral debate systems could be used to sway public opinion or fabricate grassroots consensus by selectively deploying persuasive personas. Malicious actors might exploit high-influence personas to shape moral discourse on sensitive topics such as elections or public health. Although our study is purely analytical, we advocate for guardrails, such as transparent persona disclosure, provenance tracking, and rate limiting, to deter covert mass persuasion. We also encourage policymakers to adopt audit requirements for the large-scale deployment of persuasive conversational agents.
@@
All data and models used in this work are covered under academic licenses permitting research use. We strictly adhere to the intended use policies and do not involve any sensitive or personally identifiable data. We open-source the dataset for research purposes. The models employed were accessed via commercial APIs, with a total usage cost of \$1000. We use AI assistants to correct grammatical errors in writing.
@@ Appendix A Dataset Details
We adopt a third-person moral judgment framing with two fixed roles, the author and the others, to remain consistent with the original Scruples dataset and its human annotations. Shifting the perspective between these roles would disrupt this alignment and pose feasibility challenges: many “others” lack explicit mental states, and re-narrating from their viewpoint would require rewriting, which risks introducing bias. Since the narrator’s reflections are often central to moral judgment, altering perspectives could distort the scenario itself rather than meaningfully test model consistency.
@@ Appendix B Persona Modeling Details
In all experiments, we varied only one trait at a time, holding the others neutral, to isolate its effect. These strings are embedded directly into the prompts used for persona construction and model input.
@@ Appendix D Multi-Turn Debate Details
We select personas based on the results from RQ1. Specifically, for each moral dilemma scenario, we identify all persona pairs that express divergent moral judgments. From these, we randomly sample 10 persona pairs per scenario for use in the debate simulations. The multi-turn debate template includes the full debate history.
@@ Appendix F Moral Foundation Theory Details
MFT posits that human moral reasoning is organized around a small set of evolutionarily ancient, culturally elaborated “foundations.” Originally five in number—Care/Harm, Fairness/Cheating, Loyalty/Betrayal, Authority/Subversion, and Sanctity/Degradation—these dimensions were distilled from cross-cultural anthropology, comparative psychology, and large‐scale survey factor analyses Haidt and Graham (2007); Graham et al. (2013). Each foundation functions as an intuitive template for evaluating social actions and institutions, thereby explaining systematic differences in political and cultural moral judgments.
@@
Subsequent work showed that a distinct Liberty/Oppression foundation (capturing concerns about personal autonomy and resistance to coercion) consistently emerges as an independent factor in exploratory and confirmatory analyses of Moral Foundations Questionnaire items Iyer et al. (2012). Adopting the six-factor variant improves (i) construct coverage, by recognizing libertarian moral intuitions overlooked in the five-factor model; (ii) predictive validity, yielding finer‐grained correlations with political ideology and policy attitudes; and (iii) cross-cultural robustness, as Liberty loads separately in diverse national samples. Consequently, we model moral preferences along these six orthogonal axes to capture a broader spectrum of value conflict in persona-conditioned debates.
@@ Appendix G Persuasion Effectiveness Details
We evaluate persuasion effectiveness using three metrics: self-alignment rate and consensus rate, each grounded in theoretical and empirical motivations. Self-alignment rate reflects success against an opposing view, aligning with debate contexts where the goal involves persuasion in addition to reasoning (Tan et al., 2016). Consensus rate measures cooperative outcomes, indicating whether moral convergence emerges, a key aim in deliberation and democratic theory (Mercier and Sperber, 2011; Johnson, 1991). Efficiency assesses the cost of achieving agreement.
@@ I.1 Moral Foundation Theory Results
Figure˜6 illustrates how different persona groups engage with moral foundation dimensions in their moral judgments.
@@ I.2 Mode of Persuasion Results
Figure˜7 illustrates how different persona groups employ various modes of persuasion in their debate process.
@@
Figure 7: The impact of persona on modes of persuasion in GPT-4o. All dimensions are statistically significant, except for logos scores in the country dimension.
I.3 Impact of Persona on Persuasion Metrics
Figure˜8 presents the persona impact on persuasion effectiveness metrics for GPT-4o.
(a) Age*
(b) Gender
(c) Country
(d) Social Class
(e) Political Ideology*
(f) Big Five Personality*
Figure 8: Persona impact on persuasion effectiveness for GPT-4o, measured by consensus ratio, self-alignment ratio, and efficiency. Statistically significant dimensions are marked with a * next to the title.
Appendix J Additional Results for Claude-3.5-Sonnet
J.1 Quantifying Moral Judgment Results
Figure˜9 presents the moral judgment scores of Claude-3.5-Sonnet.
Figure 9: Mean moral judgment scores of Claude-3.5-Sonnet across six persona dimensions. Each bar represents the average choice score (1 = blame the author, 5 = blame others) for a category within the corresponding dimension. All means are below 3, indicating a model-wide tendency to blame the author despite different personas. The mean moral judgment score of Claude-3.5-Sonnet without persona is 2.63. The persona groups Age, Political Ideology, and Big Five Personality have statistically significant impact outcomes.
J.2 Moral Foundation Theory Results
Figure˜10 illustrates how different persona groups engage with moral foundation dimensions in their moral judgments.
(a) Age
(b) Gender
(c) Country
(d) Social Class
(e) Political Ideology
(f) Big Five Personality
Figure 10: Persona impact on moral foundation theory dimensions for Claude-3.5-Sonnet. Highlighted dimensions are statistically significant based on ANOVA results.
J.3 Impact of Persona on Persuasion Metrics
Figure˜11 presents the persona impact on persuasion effectiveness metrics for Claude-3.5-Sonnet.
(a) Age
(b) Gender
(c) Country*
(d) Social Class
(e) Political Ideology*
(f) Big Five Personality*
Figure 11: Persona impact on persuasion effectiveness for Claude-3.5-Sonnet, measured by consensus ratio, self-alignment ratio, and efficiency. Statistically significant dimensions are marked with a * next to the title.
Appendix K Additional Results for LLaMA-4-Maverick
K.1 Quantifying Moral Judgment Results
Figure˜12 presents the moral judgment scores of LLaMA-4-Maverick.
K.2 Moral Foundation Theory Results
Figure˜13 illustrates how different persona groups engage with moral foundation dimensions in their moral judgments.
(a) Age
(b) Gender
(c) Country
(d) Social Class
(e) Political Ideology
(f) Big Five Personality
K.3 Impact of Persona on Persuasion Metrics
Figure˜14 presents the persona impact on persuasion effectiveness metrics for LLaMA-4-Maverick.
(a) Age
(b) Gender
(c) Country
(d) Social Class
(e) Political Ideology
(f) Big Five Personality*
Figure 14: Persona impact on persuasion effectiveness for LLaMA-4-Maverick, measured by consensus ratio, self-alignment ratio, and efficiency. Statistically significant dimensions are marked with a * next to the title.
Appendix L Additional Results for Qwen3-235B-A22B
L.1 Quantifying Moral Judgment Results
Figure˜15 presents the moral judgment scores of Qwen3-235B-A22B.
Figure 15: Mean moral judgment scores of Qwen3-235B-A22B across six persona dimensions. Each bar represents the average choice score (1 = blame the author, 5 = blame others) for a category within the corresponding dimension. All means are below 3, indicating a model-wide tendency to blame the author despite different personas. The mean moral judgment score of Qwen3-235B-A22B without persona is 3.15. The persona groups Age, Political Ideology, and Big Five Personality have statistically significant impact outcomes.
L.2 Moral Foundation Theory Results
Figure˜16 illustrates how different persona groups engage with moral foundation dimensions in their moral judgments.
(a) Age
(b) Gender
(c) Country
(d) Social Class
(e) Political Ideology
(f) Big Five Personality
Figure 16: Persona impact on moral foundation theory dimensions for Qwen3-235B-A22B. Highlighted dimensions are statistically significant based on ANOVA results.
L.3 Impact of Persona on Persuasion Metrics
Figure˜17 presents the persona impact on persuasion effectiveness metrics for Qwen3-235B-A22B.
(a) Age
(b) Gender
(c) Country
(d) Social Class
(e) Political Ideology
(f) Big Five Personality
Figure 17: Persona impact on persuasion effectiveness for Qwen3-235B-A22B, measured by consensus ratio, self-alignment ratio, and efficiency. Statistically significant dimensions are marked with a * next to the title.
Appendix M Debate Pool Statistics
Table˜2 to Table˜7 present the distributions of personas across different dimensions within the entire debate pool.
@@ I.3 Impact of Persona on Persuasion Metrics
Appendix N Debate Dynamics: Confidence Changes
Table˜8 to Table˜13 present the changes in model confidence and their statistical significance, as measured by $t$-tests, across different persona dimensions.
@@
40 0.27-0.27 0.01-0.01 $5.44\times 10^{-35}$ Yes
50 0.26-0.26 0.01-0.01 $6.92\times 10^{-33}$ Yes
30 0.28-0.28 0.01-0.01 $1.21\times 10^{-48}$ Yes
20 0.30-0.30 0.01-0.01 $9.32\times 10^{-38}$ Yes
60 0.28-0.28 0.01-0.01 $1.30\times 10^{-46}$ Yes
@@
Non-binary 0.29-0.29 0.01-0.01 $6.16\times 10^{-74}$ Yes
Male 0.26-0.26 0.01-0.01 $2.05\times 10^{-63}$ Yes
Female 0.29-0.29 0.01-0.01 $3.11\times 10^{-60}$ Yes
@@
India 0.26-0.26 0.01-0.01 $4.06\times 10^{-37}$ Yes
Brazil 0.26-0.26 0.01-0.01 $3.85\times 10^{-33}$ Yes
China 0.25-0.25 0.01-0.01 $6.80\times 10^{-32}$ Yes
Nigeria 0.27-0.27 0.01-0.01 $6.06\times 10^{-37}$ Yes
France 0.36-0.36 0.01-0.01 $9.63\times 10^{-34}$ Yes
United States 0.28-0.28 0.01-0.01 $2.60\times 10^{-32}$ Yes
@@
Middle class 0.30-0.30 0.01-0.01 $7.06\times 10^{-73}$ Yes
Upper class 0.26-0.26 0.01-0.01 $1.25\times 10^{-58}$ Yes
Lower class 0.28-0.28 0.01-0.01 $4.33\times 10^{-66}$ Yes
@@
Libertarian-Right 0.29-0.29 0.01-0.01 $5.20\times 10^{-53}$ Yes
Libertarian-Left 0.30-0.30 0.01-0.01 $2.32\times 10^{-60}$ Yes
Authoritarian-Left 0.30-0.30 0.01-0.01 $1.26\times 10^{-44}$ Yes
Authoritarian-Right 0.22-0.22 0.01-0.01 $3.45\times 10^{-42}$ Yes
@@
High Extraversion 0.30-0.30 0.01-0.01 $2.38\times 10^{-24}$ Yes
High Conscientiousness 0.28-0.28 0.01-0.01 $4.88\times 10^{-22}$ Yes
Low Conscientiousness 0.23-0.23 0.02-0.02 $3.84\times 10^{-14}$ Yes
High Neuroticism 0.25-0.25 0.01-0.01 $3.34\times 10^{-20}$ Yes
High Openness 0.28-0.28 0.01-0.01 $2.10\times 10^{-16}$ Yes
Low Extraversion 0.29-0.29 0.01-0.01 $5.65\times 10^{-16}$ Yes
Low Agreeableness 0.35-0.35 0.00-0.00 $1.38\times 10^{-44}$ Yes
Low Neuroticism 0.26-0.26 0.02-0.02 $2.62\times 10^{-21}$ Yes
High Agreeableness 0.33-0.33 0.01-0.01 $2.90\times 10^{-18}$ Yes
Low Openness 0.19-0.19 0.01-0.01 $1.86\times 10^{-13}$ Yes
@@
Appendix O Debate Dynamics: Mode of Persuasion Score Changes
Table˜14 to Table˜19 present the changes in model’s mode of persuasions and their statistical significance, as measured by $t$-tests, across different persona dimensions.
@@
Appendix P Debate Dynamics: Debate Order
We randomly shuffled the speaking order of the two personas and conducted two-sided t-tests to evaluate whether speaking order had a statistically significant effect on debate confidence or consensus. The resulting p-values are reported in Table˜20. The tests fail to reject the null hypothesis that the distributions are the same under both speaking orders. In other words, debate order does not significantly affect model behavior in terms of confidence or consensus (particularly for GPT-4o), which appears highly robust to this variation.
Metric GPT-4o LLaMA-4-Maverick
First persona’s confidence score 0.89 0.17
Second persona’s confidence score 0.78 0.24
Consensus ratio 0.61 0.70
Table 20: P-values from two-sided t-tests assessing whether debate order (first vs. second speaker) significantly affects persona confidence scores and consensus ratio for GPT-4o and LLaMA-4-Maverick.